# Probabilistic modeling of Cherenkov emission from particle showers

### v1

@@ 1 Introduction
Two main characteristics of Cherenkov emission from a high-energy particle shower are the total amount of emission and the profile of this emission along the shower axis. The number of Cherenkov photons emitted by a charged particle travelling at a velocity $v$ greater than the critical velocity $c/n$, per unit length, $\ell$, and wavelength, $\lambda$, is given by the Frank-Tamm formula Frank:1937fk
@@
is defined as the distance a charged particle travelling at the speed of light would need to traverse to emit the same amount of Cherenkov light as particle $i$ travelling at velocity $v_{i}$ over a distance $\ell_{i}$. Summing over $i$, the total weighted track length of all charged particles, $\hat{\ell}_{\text{tot}}$, can then be used as a measure of the total Cherenkov light emitted in a particle shower Radel:2012ij; Raedel2012. Furthermore, a gamma distribution with amplitude $\hat{\ell}_{\text{tot}}$ can be used to parametrize the amount of Cherenkov light as a function of distance along the shower axis,
@@ 2 Simulation
The first step is to produce a relatively large scale MC simulation of particle showers in ice. We use the FLUKA MC transport code Ferrari:2005zk; Ballarini:2024isa version 2025.1.0 to histogram the weighted track length of particle showers initiated by electrons, gammas, and several types of hadrons. Such MC simulations are accurate but come at a significant computational expense that increases with the initial particle energy. By generating a set of such simulations and performing a parametric model fit to each, we aim to obtain fast approximations to the overall Cherenkov photon yield and its shape.
@@ 2.1 FLUKA configuration
FLUKA is a tool for simulating particle interactions with matter and their transport, capable of reaching PeV energies for hadrons with the DPMJET hadronic interaction model Ballarini:2024isa. Configuration involves specifying the properties of the initial particle, activating relevant physical effects, describing the physical medium, and selecting appropriate scoring methods and geometry.
To provide a primary-particle-specific parametrization, simulations were conducted with a fixed set of primary particles, chosen to represent the most common final-state hadrons in neutrino DIS interactions, as well as $e^{-}$ and $\gamma$. The simulated hadron primaries include $p,\ n,\ \pi^{+},\ K^{+},\ K^{0}_{L},\ K^{0}_{S},\ \Lambda^{0},\ \Sigma^{+},\ \Sigma^{-},\ \Xi^{0},\ \Xi^{-}$, and $\Omega^{-}$. Primary particles were individually injected into a medium of ice with density $\rho_{0}=$0.9216\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{3}$\$, corresponding to that at the center of IceCube IceCube:2013llx. The initial kinetic energies of hadrons ($e^{-}$ and $\gamma$) were set to 51 (61) logarithmically spaced values from $10\text{\,}\mathrm{GeV}$ ($1\text{\,}\mathrm{GeV}$) to $1\text{\,}\mathrm{PeV}$, each with $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}\mathrm{0}$ repeated injections of a single primary particle. Each injection was assigned a specified random seed to ensure reproducibility and enable further investigation. Besides muons, all secondary particles were propagated and allowed to interact until they fell below Cherenkov-threshold energies, corresponding to $264\text{\,}\mathrm{keV}$ for electrons, positrons, and photons. Default thresholds were in place for hadrons (e.g., $100\text{\,}\mathrm{keV}$ for charged hadrons). Any muons produced were stopped immediately by setting its threshold to an impossibly high energy. These secondary muons are most commonly produced in decays of hadrons at lower energies, and a separate treatment is typically employed Panknin:2009.
The FLUKA DEFAULTS were set to PRECISIOn to turn on transport of electrons, positrons and photons and enable more precise treatment for a variety of physics processes Ballarini:2024isa. Notable additional effects that were activated via the PHYSICS command include a new evaporation model with heavy fragment evaporation, coalescence, ion splitting into nucleons, photonuclear and electronuclear interactions, and full transportation of all light and heavy ions. These settings were chosen following guidance from the FLUKA developers, and are listed in detail in appendix˜A.
@@ 2.2 Characteristics of FLUKA results
Figure 1: The left panel shows four randomly chosen $\text{d}\hat{\ell}/\text{d}x$ distributions for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ from FLUKA in solid, colored lines. The parametrization for $\text{d}\hat{\ell}/\text{d}x$ from ref. Radel:2012ij is shown as the dashed, black curve. The right panel shows the $\hat{\ell}_{\text{tot}}$ distribution for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$) from FLUKA in blue (orange). The corresponding $\hat{\ell}_{\text{tot}}$ distribution as parametrized in refs. Raedel2012; Radel:2012ij is shown as a dashed (dotted) black curves. Note that the parametrized $\hat{\ell}_{\text{tot}}$ curve for the $1\text{\,}\mathrm{TeV}$ $e^{-}$ has a very narrow width, and its mean is assumed for the $\text{d}\hat{\ell}/\text{d}x$ parametrization in the left panel.
While the average of many particle showers initiated by the same primary at a particular energy may be well approximated by a single gamma distribution Radel:2012ij, each individual shower exhibits deviations from the average. The left panel of figure˜1 shows four example $\text{d}\hat{\ell}/\text{d}x$ distributions obtained from independent FLUKA runs, in which a $1\text{\,}\mathrm{TeV}$ $e^{-}$ was injected, as solid, colored lines. For comparison, the average $\text{d}\hat{\ell}/\text{d}x$ distribution, parametrized in ref. Radel:2012ij, is shown as the dashed, black curve for the same primary. The four FLUKA runs were randomly chosen in this example, and such a comparison already demonstrates how the shower-to-shower $\text{d}\hat{\ell}/\text{d}x$ deviates from the mean, in both shape and amplitude.
The right panel of figure˜1 shows the $\hat{\ell}_{\text{tot}}$ distribution for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$) from FLUKA in blue (orange). For each run, $\hat{\ell}_{\text{tot}}$ is simply the integral of $\text{d}\hat{\ell}/\text{d}x$ over $x$, and the panel thus shows a distribution over all runs. Correspondingly, the $\hat{\ell}_{\text{tot}}$ distributions as parametrized in ref. Raedel2012; Radel:2012ij are shown as dashed (dotted) black curves. Both are Gaussians, with the $\hat{\ell}_{\text{tot}}$ from $e^{-}$ exhibiting a much smaller variance than the corresponding distribution from FLUKA. The observed difference for $e^{-}$ also exists for $\gamma$-initiated showers, and is attributable to enabling photonuclear and electronuclear effects in FLUKA Cerutti:2017tac, which were not enabled for the Geant4-based parameterizations Radel:2012ij. The physics can be broadly understood as the emission of nucleons due to excitation by MeV electrons, which translates into a decrease in the number of particles above Cherenkov threshold.
@@
A more in-depth visualization of FLUKA $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ distributions can be found in the left panel of figure˜2, which consists of 100 randomly selected profiles initiated by a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$) in blue (orange). We see that the blue lines from $e^{-}$ cluster more tightly together. In contrast, the distributions from a $\pi^{+}$ initiated shower exhibit a larger spread in shape, with a few peaking at upwards of $10\text{\,}\mathrm{m}$ from the injection point. More generally, hadron-initiated showers will exhibit larger variance in both $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ and $\hat{\ell}_{\text{tot}}$ than EM ($e^{-}$- or $\gamma$-initiated) showers Raedel2012. A comparison of the $\hat{\ell}_{\text{tot}}/E$ distributions for $e^{-}$ (blue) and $\pi^{+}$ (orange) at energies of $0.01\text{\,}\mathrm{TeV}$, $1\text{\,}\mathrm{TeV}$, and $100\text{\,}\mathrm{TeV}$ is given in the right panel of figure˜2. For $e^{-}$, the mean $\hat{\ell}_{\text{tot}}$ grows linearly with its energy. For $\pi^{+}$, the $\hat{\ell}_{\text{tot}}$ distribution has a much larger variance than its EM counterpart and deviates somewhat from linearity.
@@
Finally, we calculated Spearman’s rank coefficient, $\varrho$, to assess the correlation between $\hat{\ell}_{\text{tot}}$ and the shape and rate parameters of the gamma distribution. For each simulated shower, we fit $a$ and $b$ of eq.˜3 to $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ using non-linear least squares regression, [^2] As in ref. Radel:2012ij $x$ is in units of $l_{\text{rad}}$ and $b$ unitless, meaning the shape of the physical gamma distribution can be rescaled to varying densities.. The rank coefficient results are shown in the right panels of figure˜3, with $\varrho(\hat{\ell}_{\text{tot}},a^{\prime})$ (top), $\varrho(\hat{\ell}_{\text{tot}},b^{\prime})$ (middle), and $\varrho(a^{\prime},b^{\prime})$ (bottom), where $a^{\prime}$ and $b^{\prime}$ are defined by
@@ 3 Model
While the model from ref. Radel:2012ij; Raedel2012 is unable to describe individual particle showers, each individual shape still largely resembles a gamma distribution. The parameters of this distribution can vary quite dramatically from shower to shower. We thus parametrize each shower with equation eq.˜3, and seek a model that can generate the parameters $\hat{\ell}_{\text{tot}}$, $a$, and $b$ as random variables. The $\hat{\ell}_{\text{tot}}$ parameter controls the amplitude of the shower, while $a$ and $b$ control its shape. As shown in the right panel of figure˜3, $\hat{\ell}_{\text{tot}}$ possesses only a loose correlation with the other two parameters. Thus, a simplifying assumption was made to treat $\hat{\ell}_{\text{tot}}$ independently from the other two parameters. A more complete treatment would sample showers from the full three dimension space of $(a,b,\hat{\ell}_{\text{tot}})$ as a function of primary energy, but introduces extra complexity and computational requirements. Since shape variations are dominant, a fast and functional model for particle showers can be obtained by constructing probability distributions over the factorized space of $(a,b)$ and $\hat{\ell}_{\text{tot}}$.
@@ 3.1 Shape variations with basis splines
At each energy, the samples of $(a^{\prime},b^{\prime})$ obtained from FLUKA MC were aggregated as a two-dimensional histogram with 450 by 450 equally sized bins over the unit square. Examples for $\pi^{+}$ at three different energies are shown in the top row of figure˜5. We use this histogram data to estimate $f(a^{\prime},b^{\prime};E)$ as a function of energy. The PDF is modeled with a tensor product of exponentiated, penalized B-splines following the procedure described by Eilers and Marx 10.1214/ss/1038425655. The method is a generalized linear model (GLM), with the natural logarithm as the link function to a linear estimator. This link function ensures that the PDF is nonnegative. Error in each histogram bin is treated as being drawn from a Poisson distribution. The predictor, $\eta(a^{\prime},b^{\prime},E)=\ln f(a^{\prime},b^{\prime};E)$ is a linear combination of basis splines, which in this case are chosen to be degree three. The PDF is then given by $\exp(\eta)$. If $B_{a^{\prime},i}(a^{\prime})$, $B_{b^{\prime},j}(b^{\prime})$, and $B_{E,k}(E)$ denote the $i$th, $j$th, and $k$th basis splines in each of the three dimensions, then a single basis element for our three-dimensional spline is given by their product. Our model is thus represented by the parameter array $\bm{\theta}$, such that $\eta(a^{\prime},b^{\prime},E)=\sum_{i,j,k}\theta_{ijk}B_{a^{\prime},i}(a^{\prime})B_{b^{\prime},j}(b^{\prime})B_{E,k}(E)$.
@@
In order to have tractable fitting times we made use of general linear array modeling (GLAM) with sparse arrays. GLAM fitting speeds up the typical iterative GLM fitting procedure by orders of magnitude when working with high dimensional data. The data in this case is three dimensional, and so the array of three-dimensional basis splines that are fit to the data can be represented as the Kronecker product of three arrays of one-dimensional basis splines. By making use of properties of the Kronecker product GLAM lets us bypass the computation of particularly large matrices needed in the traditional fitting procedure. More detail on GLAM fitting can be found in ref. 10.1111/j.1467-9868.2006.00543.x.

### v2

@@ 1 Introduction
Two main characteristics of Cherenkov emission from a high-energy particle shower are the total amount of emission and the profile of this emission along the shower axis. The number of Cherenkov photons emitted by a charged particle travelling at a velocity $v$ greater than the critical velocity $c/n$, per unit length, $\ell$, and wavelength, $\lambda$, is given by the Frank-Tamm formula Frank and Tamm (1937)
@@
is defined as the distance a charged particle travelling at the speed of light would need to traverse to emit the same amount of Cherenkov light as particle $i$ travelling at velocity $v_{i}$ over a distance $\ell_{i}$. Summing over $i$, the total weighted track length of all charged particles, $\hat{\ell}_{\text{tot}}$, can then be used as a measure of the total Cherenkov light emitted in a particle shower Radel and Wiebusch (2013); Rädel (2012). Furthermore, a gamma distribution with amplitude $\hat{\ell}_{\text{tot}}$ can be used to parametrize the amount of Cherenkov light as a function of distance along the shower axis,
@@ 2 Simulation
The first step is to produce a relatively large scale MC simulation of particle showers in ice. We use the FLUKA MC transport code Ferrari et al. (2005); Ballarini and others (2024) version 2025.1.0 to histogram the weighted track length of particle showers initiated by electrons, gammas, and several types of hadrons. Such MC simulations are accurate but come at a significant computational expense that increases with the initial particle energy. By generating a set of such simulations and performing a parametric model fit to each, we aim to obtain fast approximations to the overall Cherenkov photon yield and its shape.
@@ 2.1 FLUKA configuration
FLUKA is a tool for simulating particle interactions with matter and their transport, capable of reaching PeV energies for hadrons with the DPMJET hadronic interaction model Ballarini and others (2024). Configuration involves specifying the properties of the initial particle, activating relevant physical effects, describing the physical medium, and selecting appropriate scoring methods and geometry.
To provide a primary-particle-specific parametrization, simulations were conducted with a fixed set of primary particles, chosen to represent the most common final-state hadrons in neutrino DIS interactions, as well as $e^{-}$ and $\gamma$. The simulated hadron primaries include $p,\ n,\ \pi^{+},\ K^{+},\ K^{0}_{L},\ K^{0}_{S},\ \Lambda^{0},\ \Sigma^{+},\ \Sigma^{-},\ \Xi^{0},\ \Xi^{-}$, and $\Omega^{-}$. Primary particles were individually injected into a medium of ice with density $\rho_{0}=$0.9216\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{3}$\$, corresponding to that at the center of IceCube Aartsen and others (2013). The initial kinetic energies of hadrons ($e^{-}$ and $\gamma$) were set to 51 (61) logarithmically spaced values from $10\text{\,}\mathrm{GeV}$ ($1\text{\,}\mathrm{GeV}$) to $1\text{\,}\mathrm{PeV}$, each with $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}\mathrm{0}$ repeated injections of a single primary particle. Each injection was assigned a specified random seed to ensure reproducibility and enable further investigation. Besides muons, all secondary particles were propagated and allowed to interact until they fell below Cherenkov-threshold energies, corresponding to $264\text{\,}\mathrm{keV}$ for electrons, positrons, and photons. Default thresholds were in place for hadrons (e.g., $100\text{\,}\mathrm{keV}$ for charged hadrons). Any muons produced were stopped immediately by setting its threshold to an impossibly high energy. These secondary muons are most commonly produced in decays of hadrons at lower energies, and a separate treatment is typically employed Panknin et al. (2009).
The FLUKA DEFAULTS were set to PRECISIOn to turn on transport of electrons, positrons and photons and enable more precise treatment for a variety of physics processes Ballarini and others (2024). Notable additional effects that were activated via the PHYSICS command include a new evaporation model with heavy fragment evaporation, coalescence, ion splitting into nucleons, photonuclear and electronuclear interactions, and full transportation of all light and heavy ions. These settings were chosen following guidance from the FLUKA developers, and are listed in detail in appendix˜A.
@@ 2.2 Characteristics of FLUKA results
Figure 1: The left panel shows four randomly chosen $\text{d}\hat{\ell}/\text{d}x$ distributions for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ from FLUKA in solid, colored lines. The parametrization for $\text{d}\hat{\ell}/\text{d}x$ from ref. Radel and Wiebusch (2013) is shown as the dashed, black curve. The right panel shows the $\hat{\ell}_{\text{tot}}$ distribution for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$) from FLUKA in blue (orange). The corresponding $\hat{\ell}_{\text{tot}}$ distribution as parametrized in refs. Rädel (2012); Radel and Wiebusch (2013) is shown as a dashed (dotted) black curves. Note that the parametrized $\hat{\ell}_{\text{tot}}$ curve for the $1\text{\,}\mathrm{TeV}$ $e^{-}$ has a very narrow width, and its mean is assumed for the $\text{d}\hat{\ell}/\text{d}x$ parametrization in the left panel.
While the average of many particle showers initiated by the same primary at a particular energy may be well approximated by a single gamma distribution Radel and Wiebusch (2013), each individual shower exhibits deviations from the average. The left panel of figure˜1 shows four example $\text{d}\hat{\ell}/\text{d}x$ distributions obtained from independent FLUKA runs, in which a $1\text{\,}\mathrm{TeV}$ $e^{-}$ was injected, as solid, colored lines. For comparison, the average $\text{d}\hat{\ell}/\text{d}x$ distribution, parametrized in ref. Radel and Wiebusch (2013), is shown as the dashed, black curve for the same primary. The four FLUKA runs were randomly chosen in this example, and such a comparison already demonstrates how the shower-to-shower $\text{d}\hat{\ell}/\text{d}x$ deviates from the mean, in both shape and amplitude.
The right panel of figure˜1 shows the $\hat{\ell}_{\text{tot}}$ distribution for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$) from FLUKA in blue (orange). For each run, $\hat{\ell}_{\text{tot}}$ is simply the integral of $\text{d}\hat{\ell}/\text{d}x$ over $x$, and the panel thus shows a distribution over all runs. Correspondingly, the $\hat{\ell}_{\text{tot}}$ distributions as parametrized in ref. Rädel (2012); Radel and Wiebusch (2013) are shown as dashed (dotted) black curves. Both are Gaussians, with the $\hat{\ell}_{\text{tot}}$ from $e^{-}$ exhibiting a much smaller variance than the corresponding distribution from FLUKA. The observed difference for $e^{-}$ also exists for $\gamma$-initiated showers, and is attributable to enabling photonuclear and electronuclear effects in FLUKA Cerutti et al. (2017), which were not enabled for the Geant4-based parameterizations Radel and Wiebusch (2013). The physics can be broadly understood as the emission of nucleons due to excitation by MeV electrons, which translates into a decrease in the number of particles above Cherenkov threshold.
@@
A more in-depth visualization of FLUKA $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ distributions can be found in the left panel of figure˜2, which consists of 100 randomly selected profiles initiated by a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$) in blue (orange). We see that the blue lines from $e^{-}$ cluster more tightly together. In contrast, the distributions from a $\pi^{+}$ initiated shower exhibit a larger spread in shape, with a few peaking at upwards of $10\text{\,}\mathrm{m}$ from the injection point. More generally, hadron-initiated showers will exhibit larger variance in both $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ and $\hat{\ell}_{\text{tot}}$ than EM ($e^{-}$- or $\gamma$-initiated) showers Rädel (2012). A comparison of the $\hat{\ell}_{\text{tot}}/E$ distributions for $e^{-}$ (blue) and $\pi^{+}$ (orange) at energies of $0.01\text{\,}\mathrm{TeV}$, $1\text{\,}\mathrm{TeV}$, and $100\text{\,}\mathrm{TeV}$ is given in the right panel of figure˜2. For $e^{-}$, the mean $\hat{\ell}_{\text{tot}}$ grows linearly with its energy. For $\pi^{+}$, the $\hat{\ell}_{\text{tot}}$ distribution has a much larger variance than its EM counterpart and deviates somewhat from linearity.
@@
Finally, we calculated Spearman’s rank coefficient, $\varrho$, to assess the correlation between $\hat{\ell}_{\text{tot}}$ and the shape and rate parameters of the gamma distribution. For each simulated shower, we fit $a$ and $b$ of eq.˜3 to $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ using non-linear least squares regression, [^2] As in ref. Radel and Wiebusch (2013) $x$ is in units of $l_{\text{rad}}$ and $b$ unitless, meaning the shape of the physical gamma distribution can be rescaled to varying densities.. The rank coefficient results are shown in the right panels of figure˜3, with $\varrho(\hat{\ell}_{\text{tot}},a^{\prime})$ (top), $\varrho(\hat{\ell}_{\text{tot}},b^{\prime})$ (middle), and $\varrho(a^{\prime},b^{\prime})$ (bottom), where $a^{\prime}$ and $b^{\prime}$ are defined by
@@ 3 Model
While the model from ref. Radel and Wiebusch (2013); Rädel (2012) is unable to describe individual particle showers, each individual shape still largely resembles a gamma distribution. The parameters of this distribution can vary quite dramatically from shower to shower. We thus parametrize each shower with equation eq.˜3, and seek a model that can generate the parameters $\hat{\ell}_{\text{tot}}$, $a$, and $b$ as random variables. The $\hat{\ell}_{\text{tot}}$ parameter controls the amplitude of the shower, while $a$ and $b$ control its shape. As shown in the right panel of figure˜3, $\hat{\ell}_{\text{tot}}$ possesses only a loose correlation with the other two parameters. Thus, a simplifying assumption was made to treat $\hat{\ell}_{\text{tot}}$ independently from the other two parameters. A more complete treatment would sample showers from the full three dimension space of $(a,b,\hat{\ell}_{\text{tot}})$ as a function of primary energy, but introduces extra complexity and computational requirements. Since shape variations are dominant, a fast and functional model for particle showers can be obtained by constructing probability distributions over the factorized space of $(a,b)$ and $\hat{\ell}_{\text{tot}}$.
@@ 3.1 Shape variations with basis splines
At each energy, the samples of $(a^{\prime},b^{\prime})$ obtained from FLUKA MC were aggregated as a two-dimensional histogram with 450 by 450 equally sized bins over the unit square. Examples for $\pi^{+}$ at three different energies are shown in the top row of figure˜5. We use this histogram data to estimate $f(a^{\prime},b^{\prime};E)$ as a function of energy. The PDF is modeled with a tensor product of exponentiated, penalized B-splines following the procedure described by Eilers and Marx Eilers and Marx (1996). The method is a generalized linear model (GLM), with the natural logarithm as the link function to a linear estimator. This link function ensures that the PDF is nonnegative. Error in each histogram bin is treated as being drawn from a Poisson distribution. The predictor, $\eta(a^{\prime},b^{\prime},E)=\ln f(a^{\prime},b^{\prime};E)$ is a linear combination of basis splines, which in this case are chosen to be degree three. The PDF is then given by $\exp(\eta)$. If $B_{a^{\prime},i}(a^{\prime})$, $B_{b^{\prime},j}(b^{\prime})$, and $B_{E,k}(E)$ denote the $i$th, $j$th, and $k$th basis splines in each of the three dimensions, then a single basis element for our three-dimensional spline is given by their product. Our model is thus represented by the parameter array $\bm{\theta}$, such that $\eta(a^{\prime},b^{\prime},E)=\sum_{i,j,k}\theta_{ijk}B_{a^{\prime},i}(a^{\prime})B_{b^{\prime},j}(b^{\prime})B_{E,k}(E)$.
@@
In order to have tractable fitting times we made use of general linear array modeling (GLAM) with sparse arrays. GLAM fitting speeds up the typical iterative GLM fitting procedure by orders of magnitude when working with high dimensional data. The data in this case is three dimensional, and so the array of three-dimensional basis splines that are fit to the data can be represented as the Kronecker product of three arrays of one-dimensional basis splines. By making use of properties of the Kronecker product GLAM lets us bypass the computation of particularly large matrices needed in the traditional fitting procedure. More detail on GLAM fitting can be found in ref. Currie et al. (2006).